# A comprehensive analysis toward the Fermi-LAT source 4FGL J1846.90227 Jets of a proto-planetary nebula producing -rays?

### v1

@@ 3.5 PN candidate IRAS 18443-0231
IRAS 18443-0231 appears as a planetary nebula candidate in the Simbad Catalog. It is located at the far kinematic distance of about 8 kpc with a systemic velocity of 110 km s-1 (see for instance Yang etĀ al. 2016; Irabor etĀ al. 2018). By analyzing the spectrum of the H2O maser from Urquhart etĀ al. (2011), we point out that this source could be a āwater-fountainā PN, which is indeed a very interesting kind of object due to the scarce of them (e.g. Boboltz & Marvel 2005; Fan etĀ al. 2024). As was mentioned in Section 1, 4FGL-DR3 and 4FGL-DR4 Fermi catalogs refer to the radio source TXS 1844025 (18:47:00.5, -02:27:52, J2000), which is indeed IRAS 184430231, as a āpossible associationā. Thus, we conclude that such a source deserves a thoroughest analysis in the context of this work.
Figureā10 presents in grey and green contours the radio continuum emission at 3 GHz towards the PN candidate IRAS 184430231. The blue-dashed contours represent the SCUBA continuum emission of cold dust at 850 m. The centimeter and submillimeter radio sources show a very good spatial correlation. Strikingly, the 3 GHz emission exhibits faint extended structure, whose morphology resembles jets arising from the central source.
Using the integrated flux densities at 1.4, 3.0, and 5.0 GHz, after convolving VLASS and CORNISH images to the lower resolution of the MAGPIS image, we estimate the radio spectra index (), obtaining a negative index value of (see Fig.ā11) for the whole source. In all cases the source size is smaller than the largest angular scale of each observation, therefore the radio flux should be completely recovered for the three frequencies. The background subtraction was about 5% in the MAGPIS image and negligible in the VLASS and CORNISH images. Thus, we found non-thermal radio continuum emission associated with IRAS 184430231, which indicates the presence of synchrotron particle acceleration related to jets, strongly suggesting that the source is a proto-planetary nebula (pPN).
The 850 m emission (see blue-dashed contours in Fig.ā10) suggests that IRAS 184430231 is embedded in a molecular clump. Hence, we roughly estimate the mass of gas from the cold dust emission as usually is done using the formula:
where is the integrated flux density (we obtain 0.12 Jy), the distance to the source (assumed in 8 kpc), the dust opacity per gram of matter (we used 0.0185 cm2g-1) and is the Plank function with the dust temperature (assumed in 20 K). We obtained a mass and density about 140 Mā and 103 cm-3.
@@
Additionally, we looked for molecular outflows towards this source. Using data from the JCMT we found a spectral wing in the 12CO J=3ā2 spectrum suggesting the presence of a red-shifted outflow (see Fig.ā12). It can be noticed that if there was a blue-shifted outflow it would be masked within the 12CO component centered at 95 kms-1, which correspond to foreground molecular gas. Figureā13 displays in red contours the 12CO emission integrated between 115 and 130 kms-1 (i.e. along the spectral wing shown in Fig.ā12) and clearly shows the presence of a lobe corresponding to a red-shifted molecular outflow emanating from IRAS 18443-0231.
@@
Figure 14 shows the good spatial correlation between the radio continuum emission at 3 GHz associated with the protoplanetary nebula IRAS 184430231 and the X-ray source 4XMM J184700.4022752. Moreover, it can be appreciated a diffuse extended structure toward the southwest in the X-ray emission, which seems to accompany the direction of the red lobe of the 12CO (3ā2) molecular outflow.
3.5.1 X-ray analysis of IRAS 18443-0231
@@
4XMM J184700.4022752 is a hard X-ray source, with most of the photons detected between 2.0 and . The reported count rate in the pn cameras are and for Obs1 and Obs2, respectively, indicating a constant flux between observations (refer to the 4XMM-DR13 catalog). To analyze the X-ray emission from 4XMM J184700.4022752, we extracted source and background spectra from the 6 available cameras. We obtained 289 source counts, which in principle is sufficient for a spectral fitting. Taking into account the lack of variability, we combined the source spectra of the 6 cameras into a single spectrum using the epicspeccombine task of SAS. The merged spectrum was binned with a minimum of 15 counts per bin. We used Xspec to display and analyze the spectrum. We fitted the spectrum using different emission models to determine whether the emission has a thermal or non-thermal nature.
@@
Unfortunately, we could not determine the best-fit model based on the and the model parameters had large uncertainties. Then, for the purpose of estimating the X-ray flux of the source, we assume some spectral behavior. Taking into account that the radio emission of the pPN is likely produced by synchrotron radiation and has a spectral index , we consider an absorbed power law () in the X-ray band with a fixed photon index . In other words, we are assuming that the non-thermal radio emission extends to the X-ray domain without a spectral break. We fitted the spectrum on the energy band letting the absorption () and normalization () as free parameters. We obtained , and and for 19 d.o.f. In Fig.ā15 we show the merged spectrum and the best fit powerlaw model. The unabsorbed flux in the band is and the luminosity for a distance of is . This value could be compatible with the expected non-thermal X-ray luminosity in very young pPNe (Blackman etĀ al., 2001; Sahai etĀ al., 2003).
4 Discussion about the nature of the gamma-ray emission
@@
In this section, the nature of the gamma-ray source 4FGL J1846.90227 is discussed based on the previous multiwavelength characterization of the region.
@@
Kerby etĀ al. (2021) analyzed a sample of 174 unidentified Fermi sources with a single X-ray/UV/optical counterpart. Given that the majority of the sources of 4FGL catalog are pulsars and blazars, the authors developed a neural network classifier approach using gamma-ray, X-ray, and UV/optical spectral parameters to yield a descriptive classification of unassociated spectra into pulsars and blazars. The study included the Fermi-LAT source 4FGL J1846.90227, but for the reasons mentioned in Sect.ā1, the association of this gamma-ray source with a blazar was ambiguous, which encouraged us to carry out a more in-depth study of the source.
First of all, the location of 4FGL J1846.90227, close to the galactic plane (), would suggest a galactic origin for the high energy source. Then, as shown above, we discarded the blazar nature of the the X-ray sources indicated by Kerby etĀ al. (2021). It is worth noting that the X-ray emission of the most intense one, that it is related to 4XMM J184650.6022907, was modeled by an optically thin thermal plasma (see Sect.ā3.4.2). This indeed discards a blazar as a possible counterpart, given that these extra-galactic sources typically produce non-thermal X-rays via the synchrotron mechanism (Kaur etĀ al., 2019; Kerby etĀ al., 2021).
@@ 4.2.1 X-ray binaries
In most HMXBs, the compact object usually is a neutron star, which accretes material from the stellar wind of its massive companion (Liu etĀ al., 2006). High-mass stars have powerful stellar winds with mass loss rates 10-6 Māyr-1 (Smith, 2014), so direct wind accretion usually results in high accretion rates, producing X-ray luminosities in the 10 erg s-1 range (Fornasini etĀ al., 2023). The donor star in most HMXBs is either a supergiant O/B star or a Be star, both of which have higher mass loss rates than O/B main sequence stars. The HMXBs typically have bright X-ray hard spectrum, which can be adequately modeled by a power law (Haberl etĀ al., 2022; Sidoli & Paizis, 2018), however, some of them exhibit a small excess of soft X-rays whose origin is not entirely clear (van der Meer etĀ al., 2005). Therefore, taking into account that the X-ray luminosity of 4XMM J184650.602290 is about 1030 erg s-1, that its soft X-ray spectrum is well fitted by an optically thin thermal plasma model, and finally, that the optical/IR SED indicates that the donor companion is a low-mass star (an M0 sub-giant), we can rule out that the binary BD02 4739 is a HMXB.
On the other hand, LMXBs are typically less bright than HMXBs, with X-ray luminosities that can be as low as erg s-1 for quiescent systems in which little or no accretion occurs (Bahramian & Degenaar, 2023). Therefore, at least from the X-ray luminosity point of view, we cannot rule out that 4XMM J184650.6-022907 is a LMXB. Moreover, the spectral type of the donor star, a M0 sub-giant star, is consistent with LMXBs. However, X-ray emission from LMXBs is not expected to be produced by an optically thin thermal plasma, as we found for 4XMM J184650.6022907. The spectral emission of LMXBs can be composed of three main components: disk blackbody originating from optically-thick accretion disc, a comptonized component originating due to comptonization of X-rays in the corona, and eventually, a blackbody component of emission from the neutron star surface or the boundary layer (Lin etĀ al., 2009). Regarding energy spectra modeling of LMXBs, there are two different approaches: (1) a soft/thermal, and (2) a hard/comptonized components (Mitsuda etĀ al., 1989; Hasinger & van der Klis, 1989), typically fitted by blackbody and power-law models (Abdelfatah etĀ al., 2021), respectively. In summary, given the behavior of the X-ray spectrum of 4XMM J184650.6022907 it is unlikely that the X-ray emission is originated in a LMXB.
@@ 4.2.2 A symbiotic binary system
It is worth noting that the X-ray emission from 4XMM J184650.6022907 is also compatible with the so-called -type symbiotics, first introduced by Muerset etĀ al. (1997). This class is characterized by soft emission produced by photons with energies less than . According to Sokoloski etĀ al. (2016), both burning and non-burning symbiotics may belong to the -type, in which thermal X-ray emission from an optically thin plasma arises from the gas heated by colliding winds and/or collimated jets.
@@
Typical X-ray luminosities of -type symbiotics in the band that goes from to (Luna etĀ al., 2013). The luminosities of 4XMM J184650.6022907 in this energy band are and for Obs1 and Obs2, respectively, estimated with the best-fit model of Tableā3 (right column). A striking characteristic of the X-ray emission from 4XMM J184650.6022907 is its variability, namely the flux increases by factor between 2011 and 2018. A similar behavior has been observed in AG Peg, a -type symbiotic stars that presented an outburst in 2015. This star has been extensively studied in the X-ray domain since its discover by ROSAT in 1993. It was observed with Swift in 2013 and 2015-2016, and with XMM-Newton in 2017. The X-ray emission, which was modeled by an optically thin plasma, increased by a factor 2 during the active phase in 2015, with respect to the pre- and post-outburst fluxes (Zhekov & Tomov, 2018). While the X-ray emission in the pre-outburst period is likely produced by plasma heated in the colliding-wind region of the binary system, the dominant heating mechanism during the active phase is probably accretion of the wind from the cool star into the WD (Zhekov & Tomov, 2016) or, alternatively, shocks from the interaction between material ejected during the outburst and the ejecta and/or circumbinary material (Ramsay etĀ al., 2016).
@@
Regarding the radio continuum emission towards symbiotic stars, seminal studies like Seaquist etĀ al. (1993) found that most shell burning symbiotics detected with the Very Large Array (VLA) had radio flux densities on the order of mJy, consistent with free-free emission from the ionized wind of the red giant. On the other hand, Weston (2016), studying a sample of 11 non-burning symbiotics with radio continuum observations, found that about half of the sources have faint radio flux densities of about 10 m or less. Taking this into account, we searched for possible radio counterpart for the binary star BD02 4739 using the 3 GHz radio continuum Very Large Array Sky Survey (VLASS; Lacy etĀ al., 2020). Although the high sensitivity (Jy beam-1) and the high resolution () of the data, we did not find any evidence of radio continuum emission related to the source, which is consistent with a scenario of non-burning WD-symbiotic star.
@@
Summarizing, our multiwavelength analysis strongly suggests that 4XMM J184650.6022907 is related to the spectroscopic binary star BD02 4739, which would be a non-burning -type WD-symbiotic system with a sub-giant donor star. The increase of the X-ray flux detected in the 2018 XMM-Newton observation could be a consequence of an active phase transited by the system. As pointed out by Sokoloski etĀ al. (2016), in a non shell burning symbiotic, the optical emission is dominated by the red giant. This supports our hypothesis of Sec.ā3.4.1, where we consider that the fluxes used to fit the SED come from the donor star.
4.3 Is BD02 4739 responsible for the gamma-ray emission of 4FGL J1846.90227?
@@
Symbiotic stars have been confirmed as potential gamma-ray sources in the GeV band, when they experience a nova eruption event (Li etĀ al., 2020; Franckowiak etĀ al., 2018; Abdo etĀ al., 2010). Novae are thermonuclear explosions on a WD surface fueled by mass accreted from a companion star. Symbiotic novae have an evolved companion (e.g. red giant) with a dense wind as opposed to a main-sequence companion for classical novae (Ackermann etĀ al., 2014). Current physical models suggest that shocked expanding gas from the symbiotic nova shell interacts with the dense wind of the red giant, and that particles can be accelerated effectively to produce decay -rays from proton-proton interactions (Franckowiak etĀ al., 2018). The first Fermi-LAT gamma-ray detection of a nova was made in the symbiotic-like nova V407 Cyg in 2010 (Abdo etĀ al., 2010), followed by few other examples such as V1535 Sco, V1534 Sco, and V745 Sco.
@@
We constructed a light curve of the GeV emission of 4FGL J1846.90227 using data from the Incremental Fermi LAT 4th Source Catalog (4FGL-DR3), spanning 12 years of data (August 2008 - August 2020) with 1-year bins (Fig.ā16). For comparison, we also shown the light curves of novae V5668 Sgr (4FGL J1837.62904) and V1369 Cen (4FGL J1353.35910), where we have marked the date of the optical maximum. Franckowiak etĀ al. (2018) show that gamma-ray peak of novae is concurrent with the optical maximum and lasts for some tens of days. Thus, we wonder if BD02 4739 experienced a nova event.
@@
Although the source experiences an increase in the gamma photon flux between August 2019 and August 2020, it is important to note that it presents a considerably continuous gamma emission along the years, which is not compatible with the fact that classical and symbiotic novae are only found to be GeV-emitters during the outbursts. In addition, we have not found any near-infrared/optical correspondence (VISTA, ASAS, ASAS-SN, Gaia databases) that shows a nova event. It is worth noting that based on the optical magnitude of BD-02 4739, it would be very difficult that the putative nova has been missed by both professional and amateur observatories. Moreover, by considering its nearby distance (0.5 kpc), a nova event, if not missed by Fermi, it would be extremely brighter in gamma-rays than what it is shown in Fig.ā16 (upper panel); for instance see the RS Oph 2021 case at a distance of about 2.5 kpc (Cheung etĀ al., 2022).

### v2

@@ 3.5 IRAS 184430231: a planetary nebula candidate
The systemic velocity of IRAS 184430231 is 110 km s-1 (see the RMS Database Server; Lumsden etĀ al. 2013). Yang etĀ al. (2016), in their study of distances to PN based on HI absorption, determined the far distance as the most probable for our source, which correspond to a kinematic distance of about 8.2 kpc. By analysing the spectrum of the H2O maser from Urquhart etĀ al. (2011), we point out that this source could be a āwater-fountainā PN, which is indeed a very interesting kind of object due to the scarce of them (e.g. Boboltz & Marvel 2005; Fan etĀ al. 2024). On the other hand, as it was mentioned in Section 1, 4FGL-DR3 and 4FGL-DR4 Fermi catalogues refer to the radio source TXS 1844025 (18:47:00.5, 02:27:52, J2000), which is indeed IRAS 184430231, as a āpossible associationā. Thus, we conclude that such a source deserves a thoroughest analysis in the context of this work.
Figureā10-top presents in colour and black contours the radio continuum emission at 3 GHz extracted from the second epoch of VLASS observations (VLASS 2.1) towards the PN candidate IRAS 184430231. It can be noticed a conspicuous compact radio source accompanied by a weak extended emission in the form of two not completely aligned lobes, which are clearly related to the central source. Figureā10-bottom presents in colour and black contours the radio continuum emission at 5 GHz extracted from CORNISH. It can also be appreciated the conspicuous central source and interestingly, although weaker, the vestiges of the lobes detected more clearly at 3 GHz can also be seen. Given their morphology and that they appear in both images, we wonder if these structures could be jets emerging from the central object. The dominant radio continuum emission in PNe is expected to be thermal, although some processes such as jets and magnetic fields in very young PNe (protoplanetary nebulae) could provide an environment for non-thermal emission.
S1.4GHz | S5GHz | Work | Spectral index |
(mJy) | (mJy) | () | |
905.0ā | 552.4ā | Becker etĀ al. (1994) | 0.390 0.076ā |
724.19 4.18 | 349.87 0.27 | White etĀ al. (2005) | 0.570 0.003 |
* No reported errors. | |||
Assuming arbitrary and coarse errors of 5 in flux densities. |
Therefore, we investigated the nature of the radio continuum emission in terms of its spectral index. Based on the works/catalogues of Zoonematkermani etĀ al. (1990), Becker etĀ al. (1994), and White etĀ al. (2005), which give the values of the integrated flux densities at 1.4 and 5 GHz from observations carried out with the VLA radio interferometer, we determined the radio spectral index (; from S) of the source. It is worth noting that the homogeneity of the observations of each work (angular resolution and interferometer configuration) makes completely reliable the obtained spectral index values. Tableā4 shows the integrated radio flux densities at 1.4 and 5 GHz (Cols.ā1 and 2, respectively) extracted from the Vizier catalogues associated with the works listed in Col.ā3. The estimated spectral index values (presented in Col.ā4.) are consistent with non-thermal radio continuum emission associated with the synchrotron mechanism. In the case of the data from Becker etĀ al. (1994) there are not reported errors for the flux densities, neither in the Vizier catalogue nor in the paper. Hence, we assumed an arbitrary and coarse error of 5 for such flux densities to estimate a likely upper limit for the error of the corresponding .
Frequency | Telescope | Beam | Sint | Work (Catalogue) |
(GHz) | () | (mJy) | ||
1.4 | VLA (A and B conf.) | 5 | 905 (NA) | Zoonematkermani etĀ al. (1990) |
VLA (D conf.) | 45 | 800 24 | Condon etĀ al. (1998) (NVSS) | |
VLA (B and BnA conf.) | 6 | 724 4 | White etĀ al. (2005) (MAGPIS) | |
VLA (B, C, and D conf.)ā | 6 | 678 1 | Helfand etĀ al. (2006b) (MAGPIS) | |
1.6 | VLA (C conf.) | 1314 | 574.0 1.3 | Bihr etĀ al. (2016) (THOR) |
3 | VLA (B and BnA conf.) | 2.5 | 440 23 | This workāā |
5 | VLA (C and CnD conf.) | 45 | 552 (NA) | Becker etĀ al. (1994) (GPS5) |
VLA (C and CnB conf.) | 6 | 349.87 0.27 | White etĀ al. (2005) (MAGPIS) | |
VLA (B and BnA conf.) | 1.5 | 360.35 32.91 | Purcell etĀ al. (2008) (CORNISH) | |
5.8 | VLA (B conf.) | 1 | 261.52 13.08 | Dzib etĀ al. (2023) |
8.6 | Green Bank Telescope | 82 | 752 41āāā | Anderson etĀ al. (2011) |
22 | 32 m INAF-IRA Noto Radiotelescope | 54 | 250 60āāā | Leto etĀ al. (2009) |
* VLA images combined with the Effelsberg 100 m telescope ( angular resolution). | ||||
** Using the second epoch of VLASS observations (VLASS 2.1) extracted from the Canadian Astronomy Data Centre. | ||||
*** Single dish observations whose flux densities could include extended emission contribution. |
Even though the above presented results are pretty conclusive, additionally, we carried out a rough study of the spectral index based on all IRAS 184430231 radio continuum flux densities available in the literature (see Tableā5 and Fig.ā11). With the caveat that in general such flux density points come from observations that were carried out with a variety of instruments from 1990 to 2023, and possible variability, primary-beam correction, and missing flux density effects are not taken into account, we present this complementary analysis as a support of the above obtained results. To the available radio flux densities at different frequencies (Tableā5) we added the flux density at 3 GHz estimated in this work. In the cases with several flux density values from different catalogues or works for the same frequency, the average of such values was considered (red squares in Fig.ā11). It is worth noting that it is possible that the points at 8.6 and 22 GHz (green squares in Fig.ā11), obtained from single dish observations, could include extended emission not related to the source, and hence, such flux density values could be upper limits. Two linear fittings were done: one considering all points (blue line in Fig.ā11) obtaining a radio spectral index () of , and the other one, considering only the points from interferometric observations (red line in Fig.ā11), obtaining an of .
Finally, it is important to note that if we take into account only the flux densities obtained from the six sub-bands in the range 1ā2 GHz of the THOR survey (observations acquired in July 2013), the spectral index yields a value of about (Bihr etĀ al., 2016). In spite of that such spectral range is a portion of the range analysed here, this result could be indicating the preponderance of thermal emission during a certain period in IRAS 184430231. This could be explained by the existence of variability effects, indicating that the non-thermal emission could be a transitory phenomenon in this kind of objects (see Sect.ā4.5 and SuĆ”rez etĀ al. 2015).
Finding non-thermal radio continuum emission towards this type of objects is indeed an interesting and very uncommon result. The obtained negative spectral radio index very likely indicates synchrotron emission due to particle acceleration related to jets, strongly suggesting that IRAS 184430231, catalogued as a PN candidate, is in fact a protoplanetary nebula (pPN) (see Sect.ā4.5).
In order to better understand the nature of this object, we characterise the environment in which it is embedded or located. At submillimeter wavelengths, even though it is not catalogued as a source in ATLASGAL, Bolocam, or SCUBA, we observed a very faint structure at 850 m emission from data of the last bolometer (see green contours in Fig.ā13). This suggests that IRAS 184430231 is related to a weak ( Jy beam-1; 2 above rms noise) and small (slightly larger than the beam) molecular clump.
@@
Additionally, we looked for molecular outflows towards this source. Using data from the JCMT we found a spectral wing in the 12CO J=3ā2 spectrum suggesting the presence of a red-shifted outflow (see Fig.ā12). It can be noticed that if there was a blue-shifted outflow it would be masked within the 12CO component centred at 95 kmās-1. Figureā13 displays in red contours the 12CO emission integrated between 115 and 130 kmās-1 (i.e. along the spectral wing shown in Fig.ā12) and clearly shows the presence of a lobe corresponding to a red-shifted molecular outflow emanating from IRAS 18443-0231.
3.5.1 X-ray analysis of IRAS 184430231
@@
Figureā14 shows the good spatial correlation between the radio continuum emission at 3 GHz associated with IRAS 184430231 and the X-ray source 4XMM J184700.4022752. Moreover, it can be appreciated a diffuse extended structure towards the southwest in the X-ray emission, which seems to accompany the direction of the red lobe of the 12CO molecular outflow.
@@
4XMM J184700.4022752 is a hard X-ray source, with most of the photons detected between 2.0 and . The reported count rate in the EPIC-pn cameras are and for Obs1 and Obs2, respectively, indicating no significant variability between observations (refer to the 4XMM-DR13 catalogue). To analyse the X-ray emission from 4XMM J184700.4022752, we extracted source and background spectra from the 6 cameras. We obtained 289 source counts, which in principle is sufficient for a spectral fitting. Taking into account the lack of variability, we combined the source spectra of the 6 cameras into a single spectrum using the epicspeccombine task of SAS. The merged spectrum was binned with a minimum of 15 counts per bin. We used Xspec to display and analyse the spectrum. We fitted the spectrum using different emission models to determine whether the emission has a thermal or non-thermal nature.
@@
Unfortunately, all the different models we tried were not well constraint and the parameters had large uncertainties. Then, for the purpose of estimating the X-ray flux of the source, we assumed some spectral behaviour. Taking into account that the radio emission of the pPN is likely produced by synchrotron radiation and has a spectral index , we considered an absorbed power law () in the X-ray band with a fixed photon index . In other words, we assumed that the non-thermal radio emission extends to the X-ray domain without a spectral break. We fitted the spectrum on the energy band letting the column density of the absorbing material and normalisation as free parameters. We obtained , and for 19 d.o.f. In Fig.ā15 we show the merged spectrum and the best fit power law model. The unabsorbed flux in the band is and the luminosity for a distance of is . This value is compatible with the expected non-thermal X-ray luminosity in very young pPNe (Blackman etĀ al., 2001; Sahai etĀ al., 2003).
4 Discussion about the nature of the -ray emission
@@
In this section, the nature of the -ray source 4FGL J1846.90227 is discussed based on the previous multiwavelength characterisation of the region.
@@
Kerby etĀ al. (2021) analysed a sample of 174 unidentified Fermi sources with a single X-ray/UV/optical counterpart. Given that the majority of the sources of 4FGL catalogue are pulsars and blazars, the authors developed a neural network classifier approach using -ray, X-ray, and UV/optical spectral parameters to yield a descriptive classification of unassociated spectra into pulsars and blazars. The study included the Fermi-LAT source 4FGL J1846.90227, but for the reasons mentioned in Sect.ā1, the association of this -ray source with a blazar was ambiguous, which encouraged us to carry out a more in-depth study of the source.
First of all, the location of 4FGL J1846.90227, close to the galactic plane (), would suggest a galactic origin for the high energy source. Then, as shown above, we discarded the blazar nature of the the X-ray sources indicated by Kerby etĀ al. (2021). It is worth noting that the X-ray emission of the most intense one, that it is related to 4XMM J184650.6022907, was modelled by an optically thin thermal plasma (see Sect.ā3.4.2). This indeed discards a blazar as a possible counterpart, given that these extra-galactic sources typically produce non-thermal X-rays via the synchrotron mechanism (Kaur etĀ al., 2019; Kerby etĀ al., 2021).
@@ 4.2.1 X-ray binaries
In most HMXBs, the compact object usually is a neutron star, which accretes material from the stellar wind of its massive companion (Liu etĀ al., 2006). High-mass stars have powerful stellar winds with mass loss rates 10-6 Māyr-1 (Smith, 2014), so direct wind accretion usually results in high accretion rates, producing X-ray luminosities in the 10 erg s-1 range (Fornasini etĀ al., 2023). The donor star in most HMXBs is either a supergiant O/B star or a Be star, both of which have higher mass loss rates than O/B main sequence stars. The HMXBs typically have bright X-ray hard spectrum, which can be adequately modelled by a power law (Haberl etĀ al., 2022; Sidoli & Paizis, 2018), however, some of them exhibit a small excess of soft X-rays whose origin is not entirely clear (van der Meer etĀ al., 2005). Therefore, taking into account that the X-ray luminosity of 4XMM J184650.602290 is about 1030 erg s-1, that its soft X-ray spectrum is well fitted by an optically thin thermal plasma model, and finally, that the UV/optical/IR SED indicates that the donor companion is a low-mass giant star, we can rule out that the binary BD02 4739 is a HMXB.
On the other hand, LMXBs are typically dimmer than HMXBs, with X-ray luminosities that can be as low as erg s-1 for quiescent systems in which little or no accretion occurs (Bahramian & Degenaar, 2023). Therefore, at least from the X-ray luminosity point of view, we cannot rule out that 4XMM J184650.6-022907 is a LMXB. Moreover, the characteristics of the donor star (a low-mass giant star) are consistent with LMXBs. However, X-ray emission from LMXBs is not expected to be produced by an optically thin thermal plasma, as we found for 4XMM J184650.6022907. The spectral emission of LMXBs can be composed of three main components: disk blackbody originating from optically-thick accretion disc, a comptonized component originating due to comptonization of X-rays in the corona, and eventually, a blackbody component of emission from the neutron star surface or the boundary layer (Lin etĀ al., 2009). Regarding energy spectra modelling of LMXBs, there are two different approaches: (1) a soft/thermal, and (2) a hard/comptonized components (Mitsuda etĀ al., 1989; Hasinger & van der Klis, 1989), typically fitted by blackbody and power-law models (Abdelfatah etĀ al., 2021), respectively. In summary, given the behaviour of the X-ray spectrum of 4XMM J184650.6022907 it is unlikely that the X-ray emission is originated in a LMXB.
@@ 4.2.2 A symbiotic binary system
It is worth noting that the X-ray emission from 4XMM J184650.6022907 is also compatible with the so-called -type symbiotics, first introduced by Muerset etĀ al. (1997). This class is characterised by soft emission produced by photons with energies less than . According to Sokoloski etĀ al. (2016), both burning and non-burning symbiotics may belong to the -type, in which thermal X-ray emission from an optically thin plasma arises from the gas heated by colliding winds and/or collimated jets.
@@
Typical X-ray luminosities of -type symbiotics in the band that goes from to (Luna etĀ al., 2013). The luminosities of 4XMM J184650.6022907 in this energy band are and for Obs1 and Obs2, respectively, estimated with the best-fit model of Tableā3 (right column). A striking characteristic of the X-ray emission from 4XMM J184650.6022907 is its variability, namely the flux increases by factor between 2011 and 2018. A similar behaviour has been observed in AG Peg, a -type symbiotic stars that presented an outburst in 2015. This star has been extensively studied in the X-ray domain since its discover by ROSAT in 1993. It was observed with Swift in 2013 and 2015-2016, and with XMM-Newton in 2017. The X-ray emission, which was modelled by an optically thin plasma, increased by a factor 2 during the active phase in 2015, with respect to the pre- and post-outburst fluxes (Zhekov & Tomov, 2018). While the X-ray emission in the pre-outburst period is likely produced by plasma heated in the colliding-wind region of the binary system, the dominant heating mechanism during the active phase is probably accretion of the wind from the cool star into the WD (Zhekov & Tomov, 2016) or, alternatively, shocks from the interaction between material ejected during the outburst and the ejecta and/or circumbinary material (Ramsay etĀ al., 2016).
@@
Regarding the radio continuum emission towards symbiotic stars, seminal studies like Seaquist etĀ al. (1993) found that most shell burning symbiotics detected with the Very Large Array (VLA) had radio flux densities on the order of mJy, consistent with free-free emission from the ionised wind of the red giant. On the other hand, Weston (2016), studying a sample of 11 non-burning symbiotics with radio continuum observations, found that about half of the sources have faint radio flux densities of about 10 m or less. Taking this into account, we searched for possible radio counterpart for the binary star BD02 4739 using the 3 GHz radio continuum Very Large Array Sky Survey (VLASS; Lacy etĀ al., 2020). Although the high sensitivity (Jy beam-1) and the high resolution () of the data, we did not find any evidence of radio continuum emission related to the source, which is consistent with a scenario of non-burning WD-symbiotic star.
@@
Summarising, our multiwavelength analysis strongly suggests that 4XMM J184650.6022907 is related to the spectroscopic binary star BD02 4739, which would be a non-burning -type WD-symbiotic system with a giant donor star. The increase of the X-ray flux detected in the 2018 XMM-Newton observation could be a consequence of an active phase transited by the system. As pointed out by Sokoloski etĀ al. (2016), in a non shell burning symbiotic, the optical emission is dominated by the red giant. This supports our hypothesis of Sec.ā3.4.1, where we consider that the fluxes used to fit the SED come from the donor star.
4.3 Is BD02 4739 responsible for the -ray emission of 4FGL J1846.90227?
@@
Symbiotic stars have been confirmed as potential -ray sources in the GeV band, when they experience a nova eruption event (Li etĀ al., 2020; Franckowiak etĀ al., 2018; Abdo etĀ al., 2010). Novae are thermonuclear explosions on a WD surface fuelled by mass accreted from a companion star. Symbiotic novae have an evolved companion (e.g. red giant) with a dense wind as opposed to a main-sequence companion for classical novae (Ackermann etĀ al., 2014). Current physical models suggest that shocked expanding gas from the symbiotic nova shell interacts with the dense wind of the red giant, and that particles can be accelerated effectively to produce decay -rays from proton-proton interactions (Franckowiak etĀ al., 2018). The first Fermi-LAT -ray detection of a nova was made in the symbiotic-like nova V407 Cyg in 2010 (Abdo etĀ al., 2010), followed by few other examples such as V1535 Sco, V1534 Sco, and V745 Sco.
@@
We constructed a light curve of the GeV emission of 4FGL J1846.90227 using data from the Incremental Fermi LAT 4th Source Catalogue (4FGL-DR3), spanning 12 years of data (August 2008 - August 2020) with 1-year bins (Fig.ā16). For comparison, we also shown the light curves of novae V5668 Sgr (4FGL J1837.62904) and V1369 Cen (4FGL J1353.35910), where we have marked the date of the optical maximum. Franckowiak etĀ al. (2018) show that gamma-ray peak of novae is concurrent with the optical maximum and lasts for some tens of days. Thus, we wonder if BD02 4739 experienced a nova event.
@@
Although the source experiences an increase in the -rays photon flux between August 2019 and August 2020, it is important to note that it presents a considerably continuous -ray emission along the years, which is not compatible with the fact that classical and symbiotic novae are only found to be GeV-emitters during the outbursts. In addition, we have not found any near-infrared/optical correspondence (VISTA, ASAS, ASAS-SN, Gaia databases) that shows a nova event. It is worth noting that based on the optical magnitude of BD-02 4739, it would be very difficult that the putative nova has been missed by both professional and amateur observatories. Moreover, by considering its nearby distance (0.5 kpc), a nova event, if not missed by Fermi, it would be extremely brighter in -rays than what it is shown in Fig.ā16 (upper panel); for instance see the RS Oph 2021 case at a distance of about 2.5 kpc (Cheung etĀ al., 2022).